OpenStudy (anonymous):

Find all the values of θ, 0 ≤ θ ≤ 2pi, for which 4sin^2 = 3. Description of how you do it would be golden :). Cheers.

OpenStudy (anonymous):

Actually, scratch that. All i need to know is how to find what θ is for 4sin^2 = 3.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \sin^2 \theta = \Large {3 \over 4} }\) \( \color{Black}{\Rightarrow \sin \theta =\Large {\sqrt{3} \over \sqrt4} }\) Please continue from here, as I gotta go

OpenStudy (anonymous):

Tyvm! I hit a brick wall with this question... cheers.